# Chronos: Learning the Language of Time Series

### v2

@@ 5.7 Qualitative Analysis and Limitations
where $\varepsilon_{t}\sim{\mathcal{N}}(0,1)$ and $\varphi_{1},\dots,\varphi_{p}$ are the parameters of the model. We generated time series from stationary AR processes of different orders ranging from 1 to 4, and we compared the forecasts generated by Chronos-T5 (Base) against those generated by three models: (a) the ground truth AR model that was used to generate the time series; (b) an AR model with the correct order ($p$) fitted to the time series; and (c) an AutoARIMA model fitted to the time series. Figure 14 shows the results for the AR(1) and AR(4) processes, and Figure 21 (Appendix E) shows the results for AR(2) and AR(3). We observe that Chronos-T5 (Base) generates plausible forecasts across all four AR processes. The simpler AR(1) and AR(2) processes are easier for the correctly-specified AR model and AutoARIMA model to fit, resulting in a better MSE than Chronos-T5 (Base). However, with increasing complexity in AR(3) and AR(4) processes, Chronos-T5 (Base) not only outperforms the AutoARIMA model (which belongs the same family as the ground truth model) but also performs slightly better than the fitted AR model with correct order. These results highlight that Chronos models can recognize fundamental patterns present in time series data.
@@
Using a categorical distribution to encode predictions gives Chronos flexibility in producing predictive distributions of different shapes. This is shown in Figure 15, illustrating kernel density estimate (KDE) plots of token IDs sampled from a Chronos model, for the first five time steps in the forecast horizon, across three datasets. Despite the fact that cross-entropy is not distance-aware, Chronos outputs predictive distributions over a contiguous set of tokens, and with different shapes, including multi-modal ones.
@@
One limitation of Chronos comes from the proposed tokenization approach (see Section 3.1). Specifically, the tokens we select represent values in the range $[-15s,15s]$, where $s$ is the scale of the data (mean absolute value). If $s$ is very small compared to the range of values in the series, then some observations will fall out of the representable range. An example of this behaviour is with sparse series, and as shown in Figure 16a. On the other hand, very large values of $s$ compared to the variance result in loss of precision: in the original space, tokens are spaced $30s/(B-1)$ from each other, where $B$ is the number of bins (we used $B=4094$ in our experiments); values closer than that to each other may be mapped to the same token, with an apparent loss of precision. An example of this behaviour is given in Figure 16b. Improving the tokenization to overcome these edge cases is subject for future work, but the results from Section 5.5 suggest that the Chronos models performs well on real-world data despite the limitations.
@@ 6.1 Beyond Zero-shot Univariate Forecasting
In our experiments, we evaluated Chronos in a zero-shot manner for most datasets. Such a setup highlights the competitiveness of zero-shot Chronos models against task-specific baselines. We expect that both in-domain and zero-shot results could be enhanced further through fine-tuning, an avenue we briefly explored in Section 5.5.2. This can be done using any parameter-efficient fine-tuning methods such as those based on low-rank adapters (LoRA) (Hu et al., 2021; Zhang et al., 2023). Alternatively, Chronos can be calibrated for a specific task with conformal methods (Romano et al., 2019; Stankeviciute et al., 2021; Xu & Xie, 2021). Chronos is especially attractive in the context of conformal prediction since it requires no training set, so all available data can be used for calibration.
In this work, we have focused on univariate time series forecasting since it constitutes the most common of real-world time series use-cases. Nevertheless, practical forecasting tasks often involve additional information that must be taken into account. One example involves covariates, that can be either time-independent (e.g., color of the product) or time-varying (e.g., on which days the product is on sale). Another closely related problem is multivariate forecasting, where historic values of one time series (e.g., interest rates) can influence the forecast for another time series (e.g., housing prices). The number of covariates or multivariate dimensions can vary greatly across tasks, which makes it challenging to train a single model that can handle all possible combinations. A possible solution may involve training task-specific adaptors that inject the covariates into the pretrained forecasting model (Rahman et al., 2020). As another option, we can build stacking ensembles (Ting & Witten, 1997) of Chronos and other light-weight models that excel at handling covariates such as LightGBM (Ke et al., 2017).
@@ Acknowledgements
We are indebted to Stefano Soatto for challenging us to think about the fundamental question regarding language models and time series modeling, ultimately leading to the creation of the present work. We are grateful to our fellow researchers who have contributed to this work with insightful discussions and valuable feedback, including but not limited to Devamanyu Hazarika, Imry Kissos, Laurent Callot, Baris Kurt, Valentin Flunkert, David Salinas, Boran Han, Xiaoyong Jin, Luke Huan, Youngsuk Park, Gaurav Gupta, Karthick Gopalswamy, Tim Januschowski, Jan Gasthaus, Bing Xiang, Kashif Rasul, Juba Nait Saada, Matthias Karlbauer, Mononito Goswami and Gerald Woo.
@@ Appendix A Algorithms
1:Time series datasets $\{\mathcal{X}_{1},\dots,\mathcal{X}_{N_{d}}\}$, maximum time series to be mixed $K=3$, Dirichlet concentration parameter $\alpha=1.5$, and (minimum, maximum) length of the augmented time series $(l_{\min}=128,l_{\max}=2048)$.
@@
10:$[\lambda_{1},\dots,\lambda_{k}]\sim\mathrm{Dir}(\alpha)$$\triangleright$ sample mixing weights
@@
1:Kernel bank $\mathcal{K}$, maximum kernels per time series $J=5$, and length of the time series $l_{\mathrm{syn}}=1024$.
@@ Appendix E Additional Results
This section complements Section 5.5 by providing additional details to the experimental results. Table 5 reports the training time and cost of Chronos-T5 models on a p4d.24xlarge EC2 instance. Tables 6 and 7 report the raw WQL and MASE scores together with the aggregate relative score and average rank obtained by all models on the datasets in Benchmark I. Similarly, Tables 8 and 9 report these scores on Benchmark II. Figures 18 and 19 show the average ranks obtained by different models on Benchmark I and II, respectively. Figure 20 illustrates the zero-shot performance of Chronos-T5-Synth (Small), a model trained solely on synthetic data generated using KernelSynth, against various baselines.
@@
Electricity (15 Min.) 0.076 0.076 0.081 0.081 0.077 0.319 0.105 0.103 0.082 0.090 0.091 0.189 0.079 0.081 0.084 - 0.229 - 0.117 0.279
Electricity (Hourly) 0.102 0.115 0.107 0.092 0.121 0.104 0.122 0.116 0.089 0.106 0.109 0.125 0.095 0.128 0.127 0.129 0.198 0.126 0.147 0.363
Electricity (Weekly) 0.064 0.067 0.077 0.078 0.069 0.147 0.110 0.159 0.069 0.116 0.105 0.106 0.146 0.098 0.097 0.151 0.146 0.138 0.198 0.198
KDD Cup 2018 0.273 0.272 0.294 0.271 0.359 0.369 0.286 0.277 0.252 0.330 0.280 0.571 0.312 0.302 0.315 2.266 0.521 0.528 0.556 -
London Smart Meters 0.423 0.426 0.430 0.433 0.431 0.384 0.358 0.350 0.346 0.405 0.374 0.365 0.369 0.358 0.357 - 0.660 - 0.541 0.731
M4 (Daily) 0.021 0.021 0.021 0.021 0.020 0.043 0.023 0.023 0.023 0.023 0.023 0.023 0.024 0.022 0.022 0.027 0.024 0.023 0.028 0.028
M4 (Hourly) 0.025 0.028 0.026 0.025 0.038 0.111 0.025 0.022 0.027 0.038 0.046 0.033 0.038 0.040 0.045 0.066 0.041 - 0.048 0.166
M4 (Monthly) 0.102 0.103 0.104 0.104 0.110 0.153 0.102 0.100 0.095 0.101 0.107 0.097 0.111 0.094 0.093 0.100 0.098 - 0.146 0.140
M4 (Weekly) 0.038 0.039 0.041 0.042 0.040 0.078 0.049 0.046 0.039 0.046 0.045 0.051 0.044 0.039 0.040 0.052 0.053 0.050 0.063 0.063
Pedestrian Counts 0.198 0.203 0.234 0.241 0.175 0.262 0.274 0.259 0.257 0.229 0.248 0.261 0.247 0.254 0.241 0.619 1.818 0.340 0.319 0.814
Rideshare 0.141 0.140 0.141 0.135 0.141 0.158 0.164 0.159 0.135 0.130 0.184 0.134 0.159 0.152 0.172 0.154 0.138 0.157 0.186 -
Taxi (30 Min.) 0.267 0.273 0.312 0.311 0.335 0.357 0.513 0.368 0.363 0.395 0.347 0.382 0.335 0.306 0.305 - 0.456 - 0.471 0.741
Temperature-Rain 0.663 0.670 0.686 0.704 0.669 0.717 0.655 0.685 0.804 0.718 0.708 0.670 0.848 0.780 0.798 1.182 1.060 0.869 1.424 -
Uber TLC (Daily) 0.096 0.097 0.099 0.106 0.099 0.176 0.116 0.108 0.100 0.110 0.126 0.111 0.106 0.116 0.108 0.167 0.190 0.151 0.231 0.231
Uber TLC (Hourly) 0.155 0.154 0.156 0.161 0.161 0.176 0.176 0.167 0.167 0.176 0.168 0.179 0.234 0.166 0.161 0.462 0.433 0.311 0.299 0.625
Agg. Relative Score 0.574 0.589 0.610 0.607 0.624 0.937 0.688 0.667 0.601 0.676 0.689 0.734 0.697 0.656 0.664 1.076 1.083 0.876 1.000 1.433
Avg. Rank 3.400 4.867 6.600 6.267 6.933 14.267 10.800 9.200 6.133 9.600 10.600 10.267 10.400 8.000 8.333 16.700 14.933 16.367 17.400 18.933
@@
Electricity (15 Min.) 0.403 0.409 0.430 0.453 0.406 1.169 0.706 0.626 0.450 0.515 0.637 1.108 0.452 0.579 0.567 0.508 - 0.583 - 0.498 1.270
Electricity (Hourly) 1.457 1.602 1.500 1.370 1.653 1.573 1.713 1.665 1.349 1.528 1.537 1.789 1.369 1.880 1.848 1.487 1.774 2.151 1.715 1.840 4.159
Electricity (Weekly) 1.829 1.888 2.007 2.044 1.879 2.979 2.848 2.759 1.631 2.517 1.929 2.800 2.613 1.975 2.035 1.880 3.086 3.078 3.009 3.037 3.037
KDD Cup 2018 0.663 0.673 0.688 0.656 0.771 0.844 0.661 0.656 0.616 0.779 0.671 1.022 0.695 0.674 0.731 0.737 1.014 1.138 1.023 0.994 -
London Smart Meters 0.843 0.852 0.859 0.868 0.856 0.792 0.770 0.755 0.733 0.832 0.824 0.788 0.799 0.777 0.781 0.794 - 0.966 - 0.966 1.297
M4 (Daily) 3.099 3.109 3.112 3.108 3.058 8.038 3.443 3.376 3.450 3.305 3.306 3.292 3.461 3.143 3.155 5.109 3.270 3.335 3.257 3.278 3.278
M4 (Hourly) 0.927 0.894 0.881 0.879 0.977 3.807 1.209 0.951 0.967 1.215 1.613 1.833 1.867 3.231 3.457 1.511 1.604 2.458 - 1.193 11.608
M4 (Monthly) 0.976 0.985 1.000 1.006 1.051 2.090 1.033 1.005 0.962 1.040 1.101 1.009 1.022 0.994 0.942 0.979 0.970 0.966 - 1.260 1.205
M4 (Weekly) 2.201 2.221 2.270 2.309 2.377 5.658 2.466 2.391 1.996 2.346 2.523 2.745 2.429 2.094 1.976 3.040 2.548 2.657 2.373 2.777 2.777
Pedestrian Counts 0.289 0.286 0.304 0.309 0.277 0.342 0.355 0.330 0.339 0.311 0.334 0.364 0.327 0.324 0.315 0.393 0.487 1.275 0.383 0.369 0.842
Rideshare 0.892 0.892 0.880 0.857 0.897 0.891 0.911 0.899 0.827 0.996 0.983 1.067 1.448 0.933 0.919 1.088 0.910 0.970 1.028 1.250 -
Taxi (30 Min.) 0.828 0.849 0.938 0.939 1.032 1.069 1.374 1.089 1.077 1.158 1.070 1.113 1.018 0.950 0.934 1.113 - 1.193 - 1.160 1.768
Temperature-Rain 0.982 0.991 1.013 1.033 0.984 1.031 0.963 0.988 1.250 1.015 1.076 0.994 1.370 1.232 1.343 1.226 1.968 1.945 1.524 2.243 -
Uber TLC (Daily) 0.819 0.836 0.871 0.906 0.846 1.289 0.935 0.878 0.813 0.905 0.938 0.916 0.855 0.877 0.879 0.838 1.228 1.312 1.114 1.378 1.378
Uber TLC (Hourly) 0.727 0.729 0.727 0.743 0.740 0.711 0.727 0.717 0.696 0.703 0.776 0.746 0.778 0.716 0.751 0.754 1.009 1.036 0.982 0.931 1.390
Agg. Relative Score 0.726 0.736 0.751 0.752 0.763 1.141 0.855 0.806 0.740 0.821 0.842 0.939 0.864 0.854 0.861 0.871 0.983 1.129 0.941 1.000 1.484
Avg. Rank 4.000 5.533 6.333 7.000 7.400 13.467 11.267 8.933 5.133 10.400 11.933 13.533 11.467 9.000 9.333 11.600 15.833 16.533 16.967 16.133 19.200

### v3

@@ 5.7 Qualitative Analysis and Limitations
where $\varepsilon_{t}\sim{\mathcal{N}}(0,1)$ and $\varphi_{1},\dots,\varphi_{p}$ are the parameters of the model. We generated time series from stationary AR processes of different orders ranging from 1 to 4, and we compared the forecasts generated by Chronos-T5 (Base) against those generated by three models: (a) the ground truth AR model that was used to generate the time series; (b) an AR model with the correct order ($p$) fitted to the time series; and (c) an AutoARIMA model fitted to the time series. Figure 14 shows the results for the AR(1) and AR(4) processes, and Figure 21 (Appendix E) shows the results for AR(2) and AR(3). We observe that Chronos-T5 (Base) generates plausible forecasts across all four AR processes. The simpler AR(1) and AR(2) processes are easier for the correctly-specified AR model and AutoARIMA model to fit, resulting in a better MSE than Chronos-T5 (Base). However, with increasing complexity in AR(3) and AR(4) processes, Chronos-T5 (Base) not only outperforms the AutoARIMA model (which belongs the same family as the ground truth model) but also performs on par with the fitted AR model with correct order. These results highlight that Chronos models can recognize fundamental patterns present in time series data.
@@
Using a categorical distribution to encode predictions gives Chronos flexibility in producing predictive distributions of different shapes. This is shown in Figure 15, illustrating kernel density estimate (KDE) plots of token IDs sampled from a Chronos model, for the first five time steps in the forecast horizon, across three datasets. Despite the fact that cross-entropy is not distance-aware, Chronos outputs predictive distributions over a contiguous set of tokens, and with different shapes, including multi-modal ones. Although Chronos learns the topology of the space directly from the data, we hypothesize that providing explicit topological information to the model during training may expedite the process and make the model robust for tokens where fewer datapoints are available. A potential method to inject topological information into the cross-entropy loss is through a type of label smoothing — assigning non-zero probability mass to tokens (i.e., bins) in the neighborhood of the the correct token. Farebrother et al. (2024) have obtained promising results with such a distance-aware regression-via-classification objective in the context of reinforcement learning. An in-depth theoretical and empirical analysis of the regression-via-classification paradigm in the context of time series forecasting would constitute interesting future research.
@@
One limitation of Chronos comes from the proposed tokenization approach (see Section 3.1). Specifically, the tokens we select represent bin centers in the range $[-15,+15]$, which ultimately represent original time series values in the range $[-15s,15s]$, where $s$ is the scale of the time series (mean absolute value). If $s$ is very small compared to the range of values in the series, then some observations will fall out of the representable range. An example of this behaviour is with sparse series, and as shown in Figure 16a. On the other hand, very large values of $s$ compared to the variance result in loss of precision: in the original space, tokens are spaced $30s/(B-1)$ from each other, where $B$ is the number of bins (we used $B=4094$ in our experiments); values closer than that to each other may be mapped to the same token, with an apparent loss of precision. An example of this behaviour is given in Figure 16b. An inference-time heuristic solution to this problem is to preprocess the time series using an alternative normalization scheme, such as standardization, for time series with large scale and small variance. Improving the tokenization to overcome these edge cases without heuristics is subject for future work, but the results from Section 5.5 suggest that the Chronos models performs well on real-world data despite the limitations.
@@ 6.1 Beyond Zero-shot Univariate Forecasting
In our experiments, we evaluated Chronos in a zero-shot manner for most datasets. Such a setup highlights the competitiveness of zero-shot Chronos models against task-specific baselines. We expect that both in-domain and zero-shot results could be enhanced further through fine-tuning, an avenue we briefly explored in Section 5.5.2. This can be done using any parameter-efficient fine-tuning methods such as those based on low-rank adapters (LoRA) (Hu et al., 2022; Zhang et al., 2023). Alternatively, Chronos can be calibrated for a specific task with conformal methods (Romano et al., 2019; Stankeviciute et al., 2021; Xu & Xie, 2021). Chronos is especially attractive in the context of conformal prediction since it requires no training set, so all available data can be used for calibration.
In this work, we have focused on univariate forecasting of uniformly-spaced time series since it constitutes the most common of real-world time series use-cases. Nevertheless, practical forecasting tasks often involve exogenous information that must be taken into account or may require modeling of irregularly-sampled time series (Rubanova et al., 2019; Ansari et al., 2023). One example of exogenous information is covariates, that can be either time-independent (e.g., color of the product) or time-varying (e.g., on which days the product is on sale). Another closely related problem is multivariate forecasting, where historic values of one time series (e.g., interest rates) can influence the forecast for another time series (e.g., housing prices). The number of covariates or multivariate dimensions can vary greatly across tasks, which makes it challenging to train a single model that can handle all possible combinations. A possible solution may involve training task-specific adaptors that inject the covariates into the pretrained forecasting model (Rahman et al., 2020). As another option, we can build stacking ensembles (Ting & Witten, 1997) of Chronos and other light-weight models that excel at handling covariates such as LightGBM (Ke et al., 2017).
@@ Acknowledgements
We are indebted to Stefano Soatto for challenging us to think about the fundamental question regarding language models and time series modeling, ultimately leading to the creation of the present work. We are grateful to our fellow researchers who have contributed to this work with insightful discussions and valuable feedback, including but not limited to George Karypis, Huzefa Rangwala, Devamanyu Hazarika, Imry Kissos, Laurent Callot, Baris Kurt, Valentin Flunkert, David Salinas, Boran Han, Xiaoyong Jin, Luke Huan, Youngsuk Park, Gaurav Gupta, Karthick Gopalswamy, Tim Januschowski, Jan Gasthaus, Bing Xiang, Kashif Rasul, Juba Nait Saada, Matthias Karlbauer, Hugo Senetaire, Mononito Goswami and Gerald Woo.
@@ Appendix A Algorithms
1:Time series datasets $\{\mathcal{X}_{1},\dots,\mathcal{X}_{N_{d}}\}$, maximum time series to be mixed $K=3$, symmetric Dirichlet concentration parameter $\alpha=1.5$, and (minimum, maximum) length of the augmented time series $(l_{\min}=128,l_{\max}=2048)$.
@@
10:$[\lambda_{1},\dots,\lambda_{k}]\sim\mathrm{Dir}([\alpha_{1}=\alpha,\dots,\alpha_{k}=\alpha])$$\triangleright$ sample mixing weights
@@
1:Kernel bank $\mathcal{K}$ (see table 2), maximum kernels per time series $J=5$, and length of the time series $l_{\mathrm{syn}}=1024$.
@@
Kernel Formula Hyperparameters
Constant $\kappa_{\mathrm{Const}}(x,x^{\prime})=C$ $C=1$
White Noise $\kappa_{\mathrm{White}}(x,x^{\prime})=\sigma_{n}\cdot\mathbf{1}_{(x=x^{\prime})}$ $\sigma_{n}\in\{0.1,1\}$
Linear $\kappa_{\mathrm{Lin}}(x,x^{\prime})=\sigma^{2}+x\cdot x^{\prime}$ $\sigma\in\{0,1,10\}$
RBF $\kappa_{\mathrm{RBF}}(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{2l^{2}}\right)$ $l\in\{0.1,1,10\}$
Rational Quadratic $\kappa_{\mathrm{RQ}}(x,x^{\prime})=\left(1+\frac{\|x-x^{\prime}\|^{2}}{2\alpha}\right)^{-\alpha}$ $\alpha\in\{0.1,1,10\}$
Periodic $\kappa_{\mathrm{Per}}(x,x^{\prime})=\exp\left(-2\sin^{2}\left(\pi\frac{\|x-x^{\prime}\|}{p}\right)\right)$ $p\in\{24,48,96,168,336,672,$ $7,14,30,60,365,730,$ 4,26,52,6,12,40,10}4,26,52,6,12,40,10\}4 , 26 , 52 , 6 , 12 , 40 , 10 }
Table 2: The kernel bank, $\mathcal{K}$, used in KernelSynth (algorithm 2).
@@ Appendix E Additional Results
This section complements Section 5.5 by providing additional details to the experimental results. Table 6 reports the training time and cost of Chronos-T5 models on a p4d.24xlarge EC2 instance. Tables 7 and 8 report the raw WQL and MASE scores together with the aggregate relative score and average rank obtained by all models on the datasets in Benchmark I. Similarly, Tables 9 and 10 report these scores on Benchmark II. Figures 18 and 19 show the average ranks obtained by different models on Benchmark I and II, respectively. Figure 20 illustrates the zero-shot performance of Chronos-T5-Synth (Small), a model trained solely on synthetic data generated using KernelSynth, against various baselines.
@@
Electricity (15 Min.) 0.077 0.078 0.080 0.082 0.077 0.319 0.104 0.105 0.082 0.090 0.091 0.189 0.079 0.081 0.084 - - 0.229 - 0.117 0.279
Electricity (Hourly) 0.101 0.114 0.105 0.089 0.117 0.104 0.121 0.117 0.089 0.106 0.109 0.125 0.095 0.128 0.127 0.132 0.129 0.198 0.126 0.147 0.363
Electricity (Weekly) 0.059 0.062 0.073 0.067 0.062 0.147 0.117 0.166 0.069 0.116 0.105 0.106 0.146 0.098 0.097 0.168 0.151 0.146 0.138 0.198 0.198
KDD Cup 2018 0.272 0.268 0.289 0.271 0.377 0.369 0.288 0.278 0.252 0.330 0.280 0.571 0.312 0.302 0.315 7.631 2.266 0.521 0.528 0.556 -
London Smart Meters 0.423 0.428 0.431 0.436 0.431 0.384 0.358 0.350 0.346 0.405 0.374 0.365 0.369 0.358 0.357 - - 0.660 - 0.541 0.731
M4 (Daily) 0.022 0.022 0.022 0.022 0.021 0.043 0.024 0.023 0.023 0.023 0.023 0.023 0.024 0.022 0.022 0.024 0.027 0.024 0.023 0.028 0.028
M4 (Hourly) 0.022 0.024 0.024 0.025 0.033 0.111 0.025 0.022 0.027 0.038 0.046 0.033 0.038 0.040 0.045 0.044 0.066 0.041 - 0.048 0.166
M4 (Monthly) 0.101 0.103 0.103 0.103 0.110 0.153 0.102 0.100 0.095 0.101 0.107 0.097 0.111 0.094 0.093 - 0.100 0.098 - 0.146 0.140
M4 (Weekly) 0.037 0.037 0.040 0.041 0.040 0.078 0.050 0.047 0.039 0.046 0.045 0.051 0.044 0.039 0.040 0.049 0.052 0.053 0.050 0.063 0.063
Pedestrian Counts 0.187 0.204 0.237 0.236 0.173 0.262 0.272 0.259 0.257 0.229 0.248 0.261 0.247 0.254 0.241 0.354 0.619 1.818 0.340 0.319 0.814
Rideshare 0.140 0.137 0.140 0.133 0.168 0.158 0.164 0.158 0.135 0.130 0.184 0.134 0.159 0.152 0.172 0.157 0.154 0.138 0.157 0.186 -
Taxi (30 Min.) 0.268 0.274 0.312 0.313 0.337 0.357 0.512 0.368 0.363 0.395 0.347 0.382 0.335 0.306 0.305 - - 0.456 - 0.471 0.741
Temperature-Rain 0.663 0.669 0.685 0.704 0.687 0.717 0.655 0.685 0.804 0.718 0.708 0.670 0.848 0.780 0.798 0.886 1.182 1.060 0.869 1.424 -
Uber TLC (Daily) 0.096 0.097 0.100 0.105 0.097 0.176 0.114 0.107 0.100 0.110 0.126 0.111 0.106 0.116 0.108 0.162 0.167 0.190 0.151 0.231 0.231
Uber TLC (Hourly) 0.153 0.153 0.155 0.161 0.162 0.176 0.177 0.165 0.167 0.176 0.168 0.179 0.234 0.166 0.161 0.273 0.462 0.433 0.311 0.299 0.625
Agg. Relative Score 0.564 0.580 0.603 0.598 0.623 0.937 0.691 0.670 0.601 0.676 0.689 0.734 0.697 0.656 0.664 1.060 1.076 1.083 0.876 1.000 1.433
Avg. Rank 3.400 4.667 6.200 6.067 7.533 14.533 11.133 9.133 6.333 9.533 10.733 10.400 10.467 8.200 8.533 17.367 17.200 15.333 16.567 18.000 19.667
@@
Electricity (15 Min.) 0.391 0.394 0.418 0.445 0.388 1.169 0.707 0.623 0.450 0.515 0.637 1.108 0.452 0.579 0.567 0.508 - - 0.583 - 0.498 1.270
Electricity (Hourly) 1.439 1.590 1.477 1.348 1.636 1.573 1.710 1.673 1.349 1.528 1.537 1.789 1.369 1.880 1.848 1.487 1.766 1.774 2.151 1.715 1.840 4.159
Electricity (Weekly) 1.739 1.801 1.942 1.954 1.770 2.979 2.868 2.758 1.631 2.517 1.929 2.800 2.613 1.975 2.035 1.880 3.063 3.086 3.078 3.009 3.037 3.037
KDD Cup 2018 0.683 0.646 0.687 0.667 0.881 0.844 0.662 0.656 0.616 0.779 0.671 1.022 0.695 0.674 0.731 0.737 0.971 1.014 1.138 1.023 0.994 -
London Smart Meters 0.828 0.838 0.846 0.857 0.842 0.792 0.770 0.754 0.733 0.832 0.824 0.788 0.799 0.777 0.781 0.794 - - 0.966 - 0.966 1.297
M4 (Daily) 3.144 3.160 3.148 3.154 3.079 8.038 3.448 3.377 3.450 3.305 3.306 3.292 3.461 3.143 3.155 5.109 3.224 3.270 3.335 3.257 3.278 3.278
M4 (Hourly) 0.682 0.694 0.721 0.758 0.710 3.807 1.210 0.950 0.967 1.215 1.613 1.833 1.867 3.231 3.457 1.511 1.300 1.604 2.458 - 1.193 11.608
M4 (Monthly) 0.960 0.970 0.982 0.991 1.044 2.090 1.032 1.005 0.962 1.040 1.101 1.009 1.022 0.994 0.942 0.979 - 0.970 0.966 - 1.260 1.205
M4 (Weekly) 1.998 2.021 2.113 2.155 2.225 5.658 2.484 2.448 1.996 2.346 2.523 2.745 2.429 2.094 1.976 3.040 2.394 2.548 2.657 2.373 2.777 2.777
Pedestrian Counts 0.272 0.286 0.304 0.303 0.271 0.342 0.354 0.330 0.339 0.311 0.334 0.364 0.327 0.324 0.315 0.393 0.382 0.487 1.275 0.383 0.369 0.842
Rideshare 0.865 0.862 0.854 0.830 0.921 0.891 0.910 0.900 0.827 0.996 0.983 1.067 1.448 0.933 0.919 1.088 0.944 0.910 0.970 1.028 1.250 -
Taxi (30 Min.) 0.830 0.849 0.941 0.944 1.037 1.069 1.374 1.088 1.077 1.158 1.070 1.113 1.018 0.950 0.934 1.113 - - 1.193 - 1.160 1.768
Temperature-Rain 0.980 0.986 1.012 1.029 0.974 1.031 0.963 0.988 1.250 1.015 1.076 0.994 1.370 1.232 1.343 1.226 1.625 1.968 1.945 1.524 2.243 -
Uber TLC (Daily) 0.821 0.839 0.870 0.906 0.835 1.289 0.940 0.871 0.813 0.905 0.938 0.916 0.855 0.877 0.879 0.838 1.174 1.228 1.312 1.114 1.378 1.378
Uber TLC (Hourly) 0.670 0.673 0.677 0.689 0.706 0.711 0.730 0.716 0.696 0.703 0.776 0.746 0.778 0.716 0.751 0.754 0.877 1.009 1.036 0.982 0.931 1.390
Agg. Relative Score 0.695 0.706 0.727 0.732 0.741 1.141 0.857 0.806 0.740 0.821 0.842 0.939 0.864 0.854 0.861 0.871 0.940 0.983 1.129 0.941 1.000 1.484
Avg. Rank 3.333 4.733 6.067 6.467 6.933 14.200 11.533 9.467 5.733 10.867 12.133 13.933 11.800 9.667 9.400 12.133 16.500 16.667 17.333 17.567 16.667 19.867